# Amplifying hybrid entangled states and superpositions of coherent states

### v1

@@ IV Amplification of Superpositions of Coherent States
In Fig 3., we plot fidelities and gains against the state amplitude $\alpha$ in the $(\hat{a}^{\dagger})^{2}$ case. The most noticeable feature is that only two states show different behavior for fidelity and gain. we focus on the fact that the qudit-number of two states are the last two, that is, $d-2$ and $d-1$ to explain the behavior of the two states. Recall that a qudit number $k$ is mapped to $k+2$ (mod $d$) via $(\hat{a}^{\dagger})^{2}$ amplification, thus our target states are |$\ket{\mathcal{C}^{0}_{\beta,d}}$ and |$\ket{\mathcal{C}^{1}_{\beta,d}}$. For sufficiently large $\alpha$, it does not make any special behavior due to orthogonality. However, for small amplitude, $\alpha\sim 0$, since |$\ket{\mathcal{C}^{(d-2)}_{\alpha,d}}\sim\ket{d-2}$ and |$\ket{\mathcal{C}^{(d-1)}_{\alpha,d}}\sim\ket{d-1}$, the $(\hat{a}^{\dagger})^{2}$ scheme maps most of their coefficients to |$\ket{d}$ and |$\ket{d+1}$, which do not have large overlap with |$\ket{\mathcal{C}^{0}_{\alpha,d}}\sim\ket{0}$ and |$\ket{\mathcal{C}^{1}_{\alpha,d}}\sim\ket{1}$. This makes low fidelity for the states with qudit-number $d-2$ and $d-1$. Moreover, the target states should have large amplitude $\alpha$ to have coefficients of |$\ket{d}$ and |$\ket{d+1}$, thus the ratio between the amplitude of the target state and that of the initial state diverges, which explains the behavior of the gain graphs for small $\alpha$. As $\alpha$ increases, target states have larger coefficients of |$\ket{d},\ket{2d},\dots$ and |$\ket{d+1},\ket{2d+1},\cdots$, thus their fidelity increase and gain values decrease. Furthermore, for small amplitude $\alpha$, with the same reason of troughs appears in $\hat{a}\hat{a}^{\dagger}$ amplification, fidelity is higher for |$\ket{\mathcal{C}^{(d-2)}_{\alpha,d}}$, and thus the gain is higher for |$\ket{\mathcal{C}^{(d-1)}_{\alpha,d}}$. As $\alpha$ increases, they can be reversed as one can see in Fig 3., (e) and (f) with $\alpha\approx 2.2$. Other states except for the last two show similar behavior to Fig 2. Their troughs are deeper than those of $\hat{a}\hat{a}^{\dagger}$ state because $(\hat{a}^{\dagger})^{2}$ changes photon number more than $\hat{a}\hat{a}^{\dagger}$. This makes lower fidelity and higher gain, which is also observed for hybrid states or coherent states Park16 .
@@
Now, we will investigate quantum Fisher information of the SCSs and the amplified SCSs to compare their performance of optimal phase estimation, as we did in section III. Quantum Fisher information of SCSs and amplified SCSs under $\hat{a}\hat{a}^{\dagger}$ and $(\hat{a}^{\dagger})^{2}$ schemes are calculated to be
$\mathcal{F}_{SCS}=4\alpha^{2}\frac{\sum_{n}(\alpha^{2}\omega_{d}^{(2-k)n}+\omega_{d}^{(1-k)n})\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}{\sum_{n}\omega_{d}^{-kn}\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}-4\alpha^{4}\left(\frac{\sum_{n}\omega_{d}^{(1-k)n}\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}{\sum_{n}\omega_{d}^{-kn}\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2}$ (39)
@@
$-4\left(\frac{\sum_{n}(\alpha^{6}\omega_{d}^{3n}+5\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n})\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]\omega_{d}^{-kn}}{\sum_{n}(\alpha^{4}\omega_{d}^{2n}+3\alpha^{2}\omega_{d}^{n}+1)\omega_{d}^{-kn}\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2}$ (40)
@@
$-4\left(\frac{\sum_{n}(\alpha^{6}\omega_{d}^{3n}+8\alpha^{4}\omega_{d}^{2n}+14\alpha^{2}\omega_{d}^{n}+4)\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]\omega_{d}^{-kn}}{\sum_{n}(\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n}+2)\omega_{d}^{-kn}\mathrm{Exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2}$ (41)
@@
One can see that quantum Fisher information of $d$-dimensional SCSs is not a monotonic function with respect to amplitude $\alpha$, thus we cannot follow the argument in Sec. III and Ref. Park16 that Fisher information reflects amplification gain and noiseless property. In Fig. 4, we plotted quantum Fisher information of SCSs for $d=2,3,4$ and $5$. For $d=2$, the graphs smoothly increase with amplitude, and for $d=3$, the graph shows some fluctuation but still three graphs monotonically increase with amplitude. However, for $d=4$ and $d=5$, the fluctuation width becomes so large that they lose monotonic behavior. This shows that quantum Fisher information is not a good measure of the overall performance of noiseless amplification for high dimensions. Smooth and monotonic behavior of quantum Fisher information of SCSs only appears at $d=2$, even and odd cat states, which were the main interest of Ref. Park16 . Therefore, quantum Fisher information for $d$-dimensional SCSs should be considered as a performance of optimal phase estimation, one of the criteria to compare with other optical states since phase estimation is a popular task in quantum optics, while it is a good measure of overall performance of amplification for hybrid states and $d=2$ SCSs.
@@
Figure 4: Quantum Fisher information of SCSs for (a) d=2, (b) d=3, (c) d=4 and (d) d=5.
@@
We analyze the fluctuating behavior of quantum Fisher information of SCSs with pseudo-number property. First, observe that Fock states are invariant under phase rotation. This implies that we cannot use the Fock state for phase estimation, which is reflected in quantum Fisher information that |$\mathcal{F}_{\ket{n}}=0$ for any Fock state |$\ket{n}$. As we investigated, |$\ket{\mathcal{C}^{k}_{\alpha,d}}\sim\ket{k}$ for $\alpha\sim 0$, thus their quantum Fisher information are nearly zero. As amplitude grows, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ moves to the superposition of Fock states thus its quantum Fisher information increases. As amplitude grows further where |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ begins to concentrate to some Fock state |$\ket{d\cdot n+k}$, its quantum Fisher information begins to decrease. Further growing amplitude makes |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ be a superposition of Fock states again, thus its quantum Fisher information increases… This explains the fluctuation of quantum Fisher information graphs, and it is the same analysis as that of fidelity graphs. Therefore, in the same reason, it is easily deduced why the width of fluctuation enlarges and the amplitude, where peaks and troughs appear, increases as $d$ grows.
@@
To compare quantum Fisher information of each scheme, we plot graphs of $\mathcal{R}=\mathcal{F}_{SCS}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ for $d=2,3,4,5$ in Fig. 5, as we did for hybrid states. The graph seems to coincide with our earlier qualitative analysis. However, we point out that some regions where $\hat{a}\hat{a}^{\dagger}$ yields larger quantum Fisher information than that of $\hat{a}^{\dagger}{}^{2}$ appear as $d$ increases. For example, for |$\ket{\mathcal{C}^{0}_{\alpha,5}}$, $\mathcal{R}$ is larger than $1$ in $2.51.9<\alpha<$ 2.5 region, where $\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ is in trough. For sufficiently large amplitude, $\mathcal{R}$ converges to $1$, and the two schemes do not show a difference.
@@
Figure 5: Ratio $\mathcal{R}=\mathcal{F}_{SCS}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ for (a) d=2, (b) d=3, (c) d=4, (d) d=5.
The graphs of the ratio of quantum Fisher information of amplified SCSs over those of SCSs are plotted in Fig 6., for $d=3,4,5$ with every $k$. For small amplitude, $\mathcal{F}_{N}$ of all states under the two schemes are much higher than $1$, which means that the performance of amplification for quantum phase estimation is good. After the plateau, the graphs rapidly decrease and fluctuate for SCSs with high $d$, and reach $1$ eventually. Due to the non-monotonic behavior of graphs, an amplification scheme should be chosen considering the goals of the experimentalist and the conditions of SCSs.
@@
Figure 6: Normalized Fisher information $\mathcal{F}_{N}$ of SCSs, where (a) $d=3$, $\hat{a}\hat{a}^{\dagger}$, (b) $d=3$, $\hat{a}^{\dagger}{}^{2}$, (c) $d=4$, $\hat{a}\hat{a}^{\dagger}$, (d), $d=4$, $\hat{a}^{\dagger}{}^{2}$, (e) $d=5$, $\hat{a}\hat{a}^{\dagger}$, (f) $d=5$, $\hat{a}^{\dagger}{}^{2}$. Here, $\mathcal{F}_{N}$ is defined by $\mathcal{F}^{\hat{A}}/\mathcal{F}$, where $\hat{A}$ denotes amplification scheme, $\hat{a}\hat{a}^{\dagger}$ or $\hat{a}^{\dagger}{}^{2}$
V Conclusion

### v2

@@ IV Amplification of Superpositions of Coherent States
In Fig. 3, we plot the fidelity and the optimized gain against $\alpha$ for the $\hat{a}^{\dagger}{}^{2}$ scheme. The most noticeable feature is that only two states show unusual behaviors for the fidelity and the gain. We focus on the fact that the qudit-numbers of these two states are the last two, that is, $d-2$ and $d-1$. Recall that a qudit number $k$ is mapped to $k+2$ (mod $d$) via $\hat{a}^{\dagger}{}^{2}$ amplification, thus our target states are |$\ket{\mathcal{C}^{0}_{\beta,d}}$ and |$\ket{\mathcal{C}^{1}_{\beta,d}}$. For sufficiently large $\alpha$, it does not show any unusual behavior due to the orthogonality. However, for small amplitudes, $\alpha\sim 0$, since |$\ket{\mathcal{C}^{(d-2)}_{\alpha,d}}\sim\ket{d-2}$ and |$\ket{\mathcal{C}^{(d-1)}_{\alpha,d}}\sim\ket{d-1}$, the $\hat{a}^{\dagger}{}^{2}$ scheme maps most of their coefficients to |$\ket{d}$ and |$\ket{d+1}$, which do not have a large overlap with |$\ket{\mathcal{C}^{0}_{\alpha,d}}\sim\ket{0}$ and |$\ket{\mathcal{C}^{1}_{\alpha,d}}\sim\ket{1}$. This makes low fidelity for the states with qudit-numbers $d-2$ and $d-1$. Moreover, the target states should have large amplitude $\alpha$ to have coefficients of |$\ket{d}$ and |$\ket{d+1}$, thus the ratio between the amplitude of the target state and that of the initial state diverges, which explains the behavior of the gain graphs for small $\alpha$. As $\alpha$ increases, target states have larger coefficients of |$\ket{d},\ket{2d},\dots$ and |$\ket{d+1},\ket{2d+1},\cdots$, thus their fidelity increase and gain values decrease. Furthermore, for small amplitude $\alpha$, with the same reason of troughs appears in $\hat{a}\hat{a}^{\dagger}$ amplification, fidelity is higher for |$\ket{\mathcal{C}^{(d-2)}_{\alpha,d}}$, and thus the gain is higher for |$\ket{\mathcal{C}^{(d-1)}_{\alpha,d}}$. As $\alpha$ increases, they can be reversed as one can see in Fig. 3, (e) and (f) with $\alpha\approx 2.2$. Other states except for the last two show similar behavior to Fig. 2. Their troughs are deeper than those of $\hat{a}\hat{a}^{\dagger}$ state because $\hat{a}^{\dagger}{}^{2}$ changes photon number more than $\hat{a}\hat{a}^{\dagger}$. This makes lower fidelity and higher gain, which is also observed for HESs or coherent states Park16 .
@@
Now, we investigate the quantum Fisher information of the SCSs and the amplified SCSs to compare their performance of optimal phase estimation, as we did in section III. The quantum Fisher information of SCSs and amplified SCSs under $\hat{a}\hat{a}^{\dagger}$ and $\hat{a}^{\dagger}{}^{2}$ schemes is calculated to be
$\mathcal{F}_{SCS}=4\alpha^{2}\frac{\sum_{n}(\alpha^{2}\omega_{d}^{(2-k)n}+\omega_{d}^{(1-k)n})\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}{\sum_{n}\omega_{d}^{-kn}\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}-4\alpha^{4}\left(\frac{\sum_{n}\omega_{d}^{(1-k)n}\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}{\sum_{n}\omega_{d}^{-kn}\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2},$ (39)
@@
$-4\left(\frac{\sum_{n}(\alpha^{6}\omega_{d}^{3n}+5\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n})\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]\omega_{d}^{-kn}}{\sum_{n}(\alpha^{4}\omega_{d}^{2n}+3\alpha^{2}\omega_{d}^{n}+1)\omega_{d}^{-kn}\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2},$ (40)
@@
$-4\left(\frac{\sum_{n}(\alpha^{6}\omega_{d}^{3n}+8\alpha^{4}\omega_{d}^{2n}+14\alpha^{2}\omega_{d}^{n}+4)\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]\omega_{d}^{-kn}}{\sum_{n}(\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n}+2)\omega_{d}^{-kn}\mathrm{exp}[\alpha^{2}\omega_{d}^{n}]}\right)^{2}.$ (41)
@@
One can see that the quantum Fisher information of $d$-dimensional SCSs is not a monotonic function with respect to amplitude $\alpha$, thus we cannot follow the argument in Sec. III and Ref. Park16 that Fisher information reflects amplification gain and noiseless property. In Fig. 4, we plotted quantum Fisher information of SCSs for $d=2,3,4$ and $5$. For $d=2$, the graphs smoothly increase with amplitude, and for $d=3$, the graph shows some fluctuation but still three graphs monotonically increase with amplitude. However, for $d=4$ and $d=5$, the fluctuation width becomes so large that they lose monotonic behavior. This shows that the quantum Fisher information is not a good measure of the overall performance of noiseless amplification for high dimensions. Smooth and monotonic behavior of quantum Fisher information of SCSs only appears at $d=2$, even and odd cat states, which were the main interest of Ref. Park16 . Therefore, the quantum Fisher information for $d$-dimensional SCSs should be considered as a performance of optimal phase estimation, one of the criteria to compare with other optical states since phase estimation is a popular task in quantum optics, while it is a good measure of overall performance of amplification for hybrid states and $d=2$ SCSs.
@@
Figure 4: Quantum Fisher information of SCSs($\mathcal{F}_{SCS}$) with coherent amplitudes $\alpha$ for (a) $d=2$, (b) $d=3$, (c) $d=4$, and (d) $d=5$.
@@
We analyze the fluctuating behavior of quantum Fisher information of SCSs with pseudo-number property. First, observe that Fock states are invariant under phase rotation. This implies that we cannot use the Fock state for phase estimation, which is reflected in quantum Fisher information that |$\mathcal{F}_{\ket{n}}=0$ for any Fock state |$\ket{n}$. As we investigated, |$\ket{\mathcal{C}^{k}_{\alpha,d}}\sim\ket{k}$ for $\alpha\sim 0$, thus their quantum Fisher information are nearly zero. As amplitude grows, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ is no longer rotationally invariant thus its quantum Fisher information increases. As amplitude grows further where |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ begins to concentrate to some Fock state |$\ket{d\cdot n+k}$, its quantum Fisher information begins to decrease. Further growing amplitude makes |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ be a superposition of Fock states again, thus its quantum Fisher information increases… This explains the fluctuation of quantum Fisher information graphs, and it is the same analysis as that of fidelity graphs. Therefore, in the same reason, it is easily deduced why the width of fluctuation enlarges and the amplitude, where peaks and troughs appear, increases as $d$ grows.
@@
We visualize the ratio between amplification schemes $\mathcal{R}=\mathcal{F}_{SCS}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ for $d=2,3,4,5$ in Fig. 5, as we did for hybrid states to compare quantum Fisher information of each scheme. The graph seems to coincide with our earlier qualitative analysis. However, we point out that some regions where $\hat{a}\hat{a}^{\dagger}$ yields larger quantum Fisher information than that of $\hat{a}^{\dagger}{}^{2}$ appear as $d$ increases. For example, for |$\ket{\mathcal{C}^{0}_{\alpha,5}}$, $\mathcal{R}$ is larger than $1$ in $2.51.9<\alpha<$ 2.5 region, where $\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ is in trough. For sufficiently large amplitude, $\mathcal{R}$ converges to $1$, and the two schemes do not show a difference.
@@
Figure 5: Ratio $\mathcal{R}=\mathcal{F}_{SCS}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{SCS}^{\hat{a}^{\dagger}{}^{2}}$ against coherent state amplitude $\alpha$ for (a) $d=2$, (b) $d=3$, (c) $d=4$, and (d) $d=5$.
@@
Figure 6: Normalized Fisher information $\mathcal{F}_{N}$ of SCSs against input coherent state amplitude $\alpha$, where (a) $d=3$ with $\hat{a}\hat{a}^{\dagger}$, (b) $d=3$ with $\hat{a}^{\dagger}{}^{2}$, (c) $d=4$ with $\hat{a}\hat{a}^{\dagger}$, (d) $d=4$ with $\hat{a}^{\dagger}{}^{2}$, (e) $d=5$ with $\hat{a}\hat{a}^{\dagger}$, and (f) $d=5$ with $\hat{a}^{\dagger}{}^{2}$. Here, $\mathcal{F}_{N}^{\hat{A}}$ is defined as $\mathcal{F}^{\hat{A}}/\mathcal{F}$, where $\hat{A}$ denotes an amplification scheme, $\hat{a}\hat{a}^{\dagger}$ or $\hat{a}^{\dagger}{}^{2}$, while $\mathcal{F}$ is quantum Fisher information of the input SCSs.
The graphs of the ratio of quantum Fisher information of amplified SCSs over those of SCSs are plotted in Fig. 6, for $d=3,4,5$ with every $k$. For small amplitude, $\mathcal{F}_{N}$ of all states under the two schemes are much higher than $1$, which means that the performance of amplification for quantum phase estimation is good. After the plateau, the graphs rapidly decrease and fluctuate for SCSs with high $d$, and reach $1$ eventually. Due to the non-monotonic behavior of graphs, an amplification scheme should be chosen considering the goals of the experimentalist and the conditions of SCSs.
V Success probabilities of amplification schemes
In this section, we analyze the success probabilities of both amplification schemes on HESs and SCSs. Our analysis in the previous sections is not restricted to specific methods of implementation, while the success probabilities depend on implementation schemes. Here, in order to provide some intuition regarding the success probabilities, we consider simple linear optical setups with single-photon sources and single-photon resolving detectors that distinguish between the vacuum and a single photon.
Figure 7: Schematic diagram of (a) $\hat{a}\hat{a}^{\dagger}{}$ scheme and (b) $\hat{a}^{\dagger}{}^{2}$ scheme. (c) Wigner function plot of an input SCS (|$\ket{\mathcal{C}_{1,4}^{0}}$). The $\hat{a}\hat{a}^{\dagger}{}$ scheme realized with the BSs ($\gamma=0.01$) yields a state (d, left) |$\ket{\mathcal{C}^{0,\hat{a}\hat{a}^{\dagger}{}}_{1,4}}$ which is a good approximation of (d, right) SCS with larger amplitude, |$\ket{\mathcal{C}_{1.5,4}^{0}}$. The $\hat{a}^{\dagger}{}^{2}$ scheme of (b) also yields a state of (e, left) which is a good approximation of (e, right), |$\ket{\mathcal{C}^{0,(\hat{a}^{\dagger})^{2}}_{1,4}}$, another SCS qudit with larger amplitude and different qudit number $k$, |$\ket{\mathcal{C}_{2,4}^{2}}$.
The amplification methods are depicted in Fig. 7. Let us first consider the $\hat{a}\hat{a}^{\dagger}{}$ scheme in Fig. 7(a). The single-photon addition may be performed using a single-photon state mixed with the input state at an almost transparent beam splitter followed by the vacuum measurement using a photodetector note1 ; Z2004 . The subsequent photon subtraction can be implemented by another beam splitter with high transmissivity to mix the input state with the vacuum and another photodetector that registers a single photon. The resulting state is then the photon-added-and-subtracted state on the input state. The $(\hat{a}^{\dagger}{})^{2}$ scheme may be performed in a similar way using two sequential photon addition processes as in Fig. 7(b).
The state after the photon subtraction or addition can be understood as an element of the consequent ensemble of the corresponding quantum channel. Exact forms of related Kraus operators can be described as
$\hat{K}_{-}|\psi\rangle_{s}$ $={}_{a}\langle 1|\hat{U}_{BS}|\psi\rangle_{s}|0\rangle_{a}=\sqrt{\gamma}(1-\gamma)^{\frac{\hat{n}}{2}}\hat{a},$ (42)
$\hat{K}_{+}|\psi\rangle_{s}$ $={}_{a}\langle 0|\hat{U}_{BS}|\psi\rangle_{s}|1\rangle_{a}=\sqrt{\gamma}(1-\gamma)^{\frac{\hat{n}+1}{2}}\hat{a}^{\dagger}{},$
where $K_{-}(K_{+})$ represents the Kraus operator corresponding to photon subtraction (addition). Here, $\gamma:=\mathrm{cos}^{2}\theta\ll 1$ is the transmissivity of the beam splitter unitary $U_{BS}:=e^{i\theta(\hat{a}_{s}^{\dagger}{}\hat{a}_{a}+\hat{a}_{s}\hat{a}_{a}^{\dagger}{})}$, and the subscripts mark operators whether they are in system mode ($s$) or ancilla mode ($a$).
For small $\gamma$, the operation is approximately $\hat{a}$ $(\hat{a}^{\dagger}{})$. The success probability of each operation is now given as
$P_{-}(|\psi\rangle)$ $=\mathrm{Tr}[\hat{K}_{-}^{\dagger}{}\hat{K}_{-}|\psi\rangle\langle\psi|],$ (43)
$P_{+}(|\psi\rangle)$ $=\mathrm{Tr}[(\hat{K}_{+})^{\dagger}{}\hat{K}_{+}|\psi\rangle\langle\psi|].$
The success probability of each amplification scheme, $\hat{a}\hat{a}^{\dagger}{}$ and $\hat{a}^{\dagger}{}^{2}$, can be analogously obtained from respective composite Kraus operators, $\hat{K}_{-}\hat{K}_{+}$ and $\hat{K}_{+}^{2}$ as
$P_{\hat{a}\hat{a}^{\dagger}{}}(|\psi\rangle)$ $=\mathrm{Tr}[(\hat{K}_{-}\hat{K}_{+})^{\dagger}{}\hat{K}_{-}\hat{K}_{+}|\psi\rangle\langle\psi|],$ (44)
$P_{\hat{a}^{\dagger}{}^{2}}(|\psi\rangle)$ $=\mathrm{Tr}[(\hat{K}_{+}^{2})^{\dagger}{}\hat{K}_{+}^{2}|\psi\rangle\langle\psi|].$
Figure 8: Success probabilities of the amplification schemes acting on HESs and SCSs using beam splitters with $\gamma=0.01$. Amplification success probabilities of HESs using (a) $\hat{a}\hat{a}^{\dagger}{}$ scheme and (b) $\hat{a}^{\dagger}{}^{2}$ scheme. Qudit number is not specified since the result is the same for all $k$. Amplification success probabilities of SCSs where (c) $d=3$ with $\hat{a}\hat{a}^{\dagger}{}$, (d) $d=3$ with $\hat{a}^{\dagger}{}^{2}$, (e) $d=4$ with $\hat{a}\hat{a}^{\dagger}{}$, and (f) $d=4$ with $\hat{a}^{\dagger}{}^{2}$.
Figure 8 shows the success probabilities of linear optical amplification schemes with $\gamma=0.01$ acting on HESs and SCSs. Data were obtained in two ways: numerical simulation of the linear optical circuit using a truncated Fock basis (for $n\leq 30$) and estimation of probabilities using the Kraus operators discussed above. The results of both schemes (numerical simulation and analytic calculation) are in agreement with each other. The HESs again show uniform behavior regardless of the qudit number $k$, since it is locally equivalent to a classical mixture of coherent states with the same amplitude.
The SCSs show more complicated behaviors. For large $\alpha$, the qudit number $k$ does not effectively affect the success probability, just as the cases for other quantities analyzed in the previous sections (fidelity, optimized gain, and QFI). For small amplitudes, however, the SCSs are approximated as Fock states as $|\mathcal{C}^{k}_{\alpha,d}\rangle\approx|k\rangle$. In this regime, the effects of the photon addition and the subtraction are clearly distinct depending on $k$, and the change of the qudit number $k$ clearly changes the average photon number of the state. This is not the case for large $\alpha$ where $|\mathcal{C}^{k}_{\alpha,d}\rangle$’s have nearly the same average photon number for given $\alpha$ regardless of the value of $k$. In fact, the leading order term of both the Kraus operators ($\hat{K}_{+}$ and $\hat{K}_{-}$) for small gamma is $\gamma\hat{n}$. Therefore, the changes of $k$ considerably affects the success probability only for small $\alpha$. The leading term of the Kraus operators also implies the uniform increase in the larger amplitude regime by increasing $\alpha$ regardless of the qudit number and the amplification scheme.
We can thus generally point out that a larger success probability can be achieved by an input state with a larger amplitude, while it has a trade-off of smaller gain. When $\alpha$ is small, it is better to choose a large value of $k$ in order to increase the success probability. High success probabilities may also be achieved by increasing the beam-splitter transmissivity $\gamma$, but it would decrease the fidelity.
VI Conclusions
We have investigated two amplification schemes, the photon-addition-and-then-subtraction scheme ($\hat{a}^{\dagger}\hat{a}$) and the successive-photon-addition scheme ($\hat{a}^{\dagger}{}^{2}$), applied to two types of nonclassical states, i.e., HESs and SCSs. A difference between HESs and SCSs is that the former is a two-mode state with entanglement while the latter is a single-mode state. A HES is locally indistinguishable from a classical mixture of two coherent states if only the coherent-state mode is considered. Thus, it is natural that the results of the amplification schemes with a HES are similar to those with a coherent state of the same amplitude. However, the SCSs are single-mode superpositions of two coherent states and their behaviors are different. In a small amplitude regime, the SCSs are similar to the Fock states with a photon number same as their qudit number. With larger amplitudes, however, the SCSs are rather closer to coherent states with the same amplitudes except for the additional modular photon number condition.